OpenStudy (anonymous):

In a sample of 200 children, 75 were found to be lovely . In a school with 100,000 children, how many would you expect to be lovely ? Plzzz HELP!!!!!*

Mehek (mehek14):

\(\dfrac{75}{200}=\dfrac{x}{100,000}\)

Mehek (mehek14):

@Halia

OpenStudy (anonymous):

Thank you so much @mehek14

Mehek (mehek14):

do you know how to solve?

OpenStudy (anonymous):

Not really I forgot 😞 I'm just here trying to catch up on this summer homework

Mehek (mehek14):

cross multiply \(75*100,000=7,500,000\\ \dfrac{7,500,000}{200}=37500\)

Mehek (mehek14):

so your answer is 37,500

OpenStudy (anonymous):

Thank you so much again I'll work it out right now

Mehek (mehek14):

ok